# Fast Subspace Fluid Simulation with Temporal-Aware Basis

### v1

@@ 6.3.1. Evolution
Results are shown on Figure 9. We see that even if the initial velocity field is significantly different from the input field, the low-order basis is able to capture the bulk flow of the low-resolution input, and modify the DMD-produced field accordingly. In particular, we note that naively applying the DMD operator, without passing the low-resolution input field into the low-order bases, ends up reconstructing the original training set, rather than a velocity field directed by our input. This is demonstrated by the results for the two initial conditions being very similar, whereas our augmented field matches the input much closer.
@@ 6.3.2. Projection
As a sanity check, we show the effect of this projection here: it is apparent with the projection step,we can recover fields that are much closer to the input, yet retaining extra high-order detail. And of course, because these are all linear, linear combinations of the direct and projected fields can be taken. In particular, because the basis functions of reduced space are orthogonal, a diagonal matrix of linear weights can be taken, preferring projected for low-order modes and direct for high-order modes for example.
@@
We also note that as a spectral decomposition of the reduced Koopman operator, much of the directability power of DMD comes from the exposure of the complex-valued eigenvalues. Because each eigenvalue represents a different speed of rotation, and their corresponding eigenvectors represent their spatial bases, we gain manipulability over different scales of turbulence for free. This is similar in concept to other spectral methods (Kim et al., 2008), but with corresponding spatial bases that learned from input data rather than predefined bases.
We thus show that DMD straddles the realm between spatial order reduction (PCA and co.) and temporal order reduction (spectral methods), gaining strengths of both methodologies.
@@ 7.1. Limitations
It is well-known that reduced-order model trades generalization for performance; our method is no exception. We have demonstrated that we can represent velocity fields that were unseen by DMD, as well as interactively push current states towards modified states. Despite this, representability is still a key limitation–there just may not be the spatial basis to support certain velocity fields. Thus, users may find that interacting with the fields, especially in states far away from observed data, behave strangely. This warrants further work on improving spatial generalization of these methods.
@@ Acknowledgements.
Appendix A Additional Results
Figure 10. Additional plume simulation results in Fig. 3.
@@ Appendix A Additional Results
Additional plume simulation results in Fig. 3. Comparison of fluid simulation results using Dynamic Mode Decomposition (DMD) with varying numbers of basis functions ($r$).** The leftmost column (MC) represents the Monte Carlo (ground truth) simulation. The remaining columns depict reconstructions using DMD with increasing numbers of basis functions ($r=2,9,28,61,105,130,150$). Rows correspond to different time steps ($t=100,150,200,250,300$). As $r$ increases, the reconstruction quality improves, capturing more details of the plume-like fluid structure over time.
Figure 11. Additional sphere simulation results in Fig. 3.
[…]
Additional sphere simulation results in Fig. 3. Comparison of fluid simulation results using Dynamic Mode Decomposition (DMD) with varying numbers of basis functions ($r$).** The leftmost column (MC) represents the Monte Carlo (ground truth) simulation. The remaining columns depict reconstructions using DMD with increasing numbers of basis functions ($r=2,9,28,61,105,130,150$). Rows correspond to different time steps ($t=100,150,200,250,300$). As $r$ increases, the reconstruction quality improves, capturing more details of the plume-like fluid structure over time.
Figure 12. Additional bunny simulation results in Fig. 3.
[…]
Additional bunny simulation results in Fig. 3. Comparison of fluid simulation results using Dynamic Mode Decomposition (DMD) with varying numbers of basis functions ($r$).** The leftmost column (MC) represents the Monte Carlo (ground truth) simulation. The remaining columns depict reconstructions using DMD with increasing numbers of basis functions ($r=2,9,28,61,105,130,150$). Rows correspond to different time steps ($t=100,150,200,250,300$). As $r$ increases, the reconstruction quality improves, capturing more details of the plume-like fluid structure over time.

### v2

@@ 6.3.1. Evolution
Results are shown on Fig. 9. We see that even if the initial velocity field is significantly different from the input field, the low-order basis is able to capture the bulk flow of the low-resolution input, and modify the DMD-produced field accordingly. In particular, we note that naively applying the DMD operator, without passing the low-resolution input field into the low-order bases, ends up reconstructing the original training set, rather than a velocity field directed by our input. This is demonstrated by the results for the two initial conditions being very similar, whereas our augmented field matches the input much closer.
@@ 6.3.2. Projection
As a sanity check, we show the effect of this projection here: it is apparent with the projection step,we can recover fields that are much closer to the input, yet retaining extra high-order detail. And, of course, because these are all linear, linear combinations of the direct and projected fields can be taken. In particular, because the basis functions of reduced space are orthogonal, a diagonal matrix of linear weights can be taken, preferring projected for low-order modes and direct for high-order modes, for example.
@@
We also note that as a spectral decomposition of the reduced Koopman operator, much of the directability power of DMD comes from the exposure of the complex-valued eigenvalues. Because each eigenvalue represents a different speed of rotation, and their corresponding eigenvectors represent their spatial bases, we gain manipulability over different scales of turbulence for free. This is similar in concept to other spectral methods (Kim et al., 2008), but with corresponding spatial bases that learned from input data rather than predefined bases. We thus show that DMD straddles the realm between spatial order reduction (PCA and co.) and temporal order reduction (spectral methods), gaining strengths of both methodologies.
The other interesting observation is that dynamic mode decomposition (DMD) shares the same algebraic backbone as a first-order vector autoregression, VAR(1) (Stock and Watson, 2001): both estimate a linear map that advances a state from $t$ to t+1t{+}1. Under a noise-free assumption, the two time-shifted snapshot matrices $X^{\prime}$ and $X$ introduced in Eq. (4) satisfy the VAR(1) relation
(18) $X^{\prime}=\bm{A}X,\quad\bm{A}\in\mathbb{R}^{(Nn)\times(Nn)},$
with $\bm{A}$ interpreted as the state-transition matrix. DMD solves the identical least-squares problem but replaces the full $\bm{A}$—the restricted Koopman operator $\bm{K}$ in our notation—with its truncated SVD, yielding a compact low-rank approximation that captures the dominant dynamical modes.
@@ 7.1. Limitations
It is well-known that reduced-order model trades generalization for performance; our method is no exception. We have demonstrated that we can represent velocity fields that were unseen by DMD, as well as interactively push current states towards modified states. Despite this, representability is still a key limitation–there just may not be the spatial basis and the reduced Koopman operator to support certain velocity fields and their nonlinear evolution. Thus, users may find that interacting with the fields, especially in states far away from observed data, behave strangely. This warrants further work on improving spatial generalization of these methods.
@@ Acknowledgements.
We are grateful to Jonathan Chalaturnyk for illuminating discussions that shaped the early stages of this research. We thank Xinwen Ding for helpful proofreading feedback. This research was made possible with the administrative support of our lab’s system administrator, John Hancock, and financial officer, Xuan Dam. This research was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through grant RGPIN-2021-03733, whose funding made this work possible.
Appendix A Koopman Theory
We briefly summarize Koopman theory, following the presentation in Brunton et al. (2022), to motivate the the linearization introduced in Eq. (3). Once again, consider the autonomous system:
(19) $\frac{d\bm{u}}{dt}=\bm{f}(\bm{u}).$
The time-dependent state $\bm{u}(t)$ subject to this dynamical system and some initial condition $\bm{u}(0)=\bm{u}_{0}$ may be represented with a time-dependent family of flow maps $\{\bm{F}_{t}\}_{t>0}:\mathbb{R}^{Nn}\to\mathbb{R}^{Nn}$ such that:
(20) $\bm{u}(t)=\bm{F}_{t}(\bm{u}_{0}).$
That is, it maps some initial state $\bm{u}_{0}$ from time 0 to its state at time $t$. In general, $\bm{F}(\bm{u})$ is not a linear map, but as per Koopman and Neumann (1932) may be lifted in a higher space that admits a linear mapping in that space. In particular, take a Hilbert space $\mathcal{H}\subseteq\{g:\mathbb{R}^{N}\to\mathbb{C}\}$ of scalar observation functions $g\in\mathcal{H}$. With these observables, Koopman (1931) defines a family of Koopman operators $\bm{\mathcal{K}}_{t}:\mathcal{H}\to\mathcal{H}$ such that
@@ Appendix A Koopman Theory
forming a trajectory in $\mathcal{H}$. That is to say, $\bm{\mathcal{K}}_{t}$ maps the measurement operator $g$ to its state at time $t$. Consequently, rather than following the nonlinear flow according to $\bm{F}_{t}$ of some initial state $\bm{u}_{0}$, we can instead take an observation of $\bm{u}_{0}$, and linearly evolve it according to $\mathcal{K}_{t}$. Notice that because $\mathcal{H}$ is a linear space, it follows from Eq. (21) that:
(22) $\mathcal{K}_{t}(\alpha g_{\alpha}+\beta g_{\beta})=\alpha\mathcal{K}_{t}g_{\alpha}+\beta\mathcal{K}_{t}g_{\beta},$
irrespective of whether $\bm{F}_{t}$ is linear in state space $\mathbb{R}^{N}$.
An infinitesimal generator $\mathcal{L}$ of the family $\{\mathcal{K}_{t}\}$ exists given $f$ is an autonomous system (Eq. (19)), and can be found via the limit,
(23) $\mathcal{L}(g(\bm{u}(t)))=\lim_{\tau\to 0}\frac{\mathcal{K}_{\tau}g(\bm{u}(t))-g(\bm{u}(t))}{\tau}.$
Notice that the limit is exactly the definition of a derivative, so that:
(24) $\frac{d}{dt}g(\bm{u}(t))=\mathcal{L}g(\bm{u}(t)).$
This is analogous to Eq. (19), but with the finite nonlinear operator $\bm{f}$ replaced by an infinite-dimensional linear operator $\mathcal{L}$. This has the matrix exponential solution $\bm{g}(\bm{u}(t+\tau))=e^{\mathcal{\mathcal{L}}(t+\tau)}\bm{g}(\bm{u}(t))$. Thus, for discrete time $t\in\mathbb{N}$, we can simply take some constant timestep $\tau$ to find a single Koopman operator $\mathcal{K}=e^{\mathcal{L}\tau}$, notably acting linearly in $\mathcal{H}$, whose $m$-times repeated application $\{\mathcal{K}^{m}=\mathcal{K}\circ\mathcal{K}\circ\cdots\circ\mathcal{K}\}$ constructs the family of operators at discrete multiples of $\tau$. This discrete-time Koopman operator, often called the Koopman operator, thus giving the timestepping scheme:
(25) $g(\bm{u}((k+1)\tau))=\mathcal{K}g(\bm{u}(k\tau)).$
Now, suppose we have some input data consisting of a (necessarily) finite set of measurements. We can consider each one of these measurements to be exactly one evaluation of the observation function. We can thus restrict our set of observation functions only to those required for the given measurements. That is, we only need to consider observation functions spanned by the set $\{g_{i}(\bm{u})=\bm{u}(\bm{x}_{i})\}$, where $\bm{x}_{i}$’s are all the spatial degrees of freedom of our measurements. In other words, we limit $\mathcal{H}$ only to the subspace of observables spanned by input data. As a convenient representation of the entire system observation function, we define the stacked vector,
(26) $\bm{g}(\bm{u})=\begin{bmatrix}g_{1}(\bm{u})\\ g_{2}(\bm{u})\\ \vdots\\ g_{n}(\bm{u})\end{bmatrix}=\begin{bmatrix}\bm{u}(\bm{x}_{1})\\ \bm{u}(\bm{x}_{2})\\ \vdots\\ \bm{u}(\bm{x}_{n})\end{bmatrix}$
where $n$ is the total spatial degrees of freedom of our measurements. We then define $\bm{K}$ to be the Koopman operator restricted to this subspace, providing a linearized timestepping operator,
(27) $\bm{g}_{k+1}=\bm{K}\bm{g}_{k},$
analogous to Eq. (25).
We have thus shown that the finite-dimensional nonlinear system provided in Eq. (19) may be lifted into a linear system in an infinite-dimensioned Hilbert space $\mathcal{H}$, then restricted to a finite linear system respecting a finite set of measurements. We would like to emphasize here, as a key limitation of the theory, the requirement of the nonlinear system to be autonomous. That is, the dynamics is necessarily only state-dependent and time-independent. This means that we cannot hope to train an operator that includes some time-dependent source term. Consequently, our moduli-editing schemes cannot be seen as solving the original system given by Eq. (19), but rather provides a perturbed version of the original solution.
Appendix B Additional Results
Figure 10. Additional plume simulation results in Fig. 3.:
Figure 11. Additional plume with sphere simulation results in Fig. 3.:
Figure 12. Additional plume with bunny simulation results in Fig. 3.: